# Dr.Mi-Bench: A Modular-integrated Benchmark for Scientific Deep Research Agent

### v1

@@ Abstract
The explosive growth in academic literature necessitates automated deep research (DR) agents, yet their evaluation remains a significant challenge. First, existing benchmarks often focus narrowly on retrieval while neglecting high-level planning and reasoning. Second, existing benchmarks favor general domains over the scientific domains that are the core application for DR agents. To address these gaps, we introduce Dr.Mi-Bench, a Modular-integrated benchmark for scientific DR agents. Grounded in academic literature, our benchmark is a human-annotated dataset of 200 instances across 10 scientific domains, including both research and review papers. Besides, we also propose a Modular-integrated Evaluation Paradigm for DR Agents (Dr.Mi-Eval), which leverages the rich structure of academic papers to assess the core competencies of planning, retrieval, and reasoning through two complementary modes: an end-to-end evaluation for DR agents and an isolated evaluation for foundational LLMs as potential backbones. Experimental results reveal a fragmented performance landscape: agents exhibit specialized strengths but share critical weaknesses, most notably in performing the multi-source retrieval required for review-style tasks and performing consistently across diverse scientific fields. Moreover, improving high-level planning capability is the crucial factor for unlocking the reasoning potential of foundational LLMs as backbones. By exposing these actionable failure modes, Dr.Mi-Bench provides a diagnostic tool to guide the development of more reliable academic research assistants.
@@ 1 Introduction
Human society has entered an era of unprecedented information explosion, making it increasingly impractical for human experts to handle the intensive knowledge work required for deep, accurate, and reliable research with conventional manual methods (zhang2025surveytesttimescalinglarge). The scale and velocity of academic publication are particularly staggering. For instance, more than 64 million research papers have become accessible online since 1996 (wordsrated2025), and the pace of new contributions continues to accelerate. This makes completing a comprehensive report with conventional manual methods increasingly difficult, thereby underscoring the urgent need for automated systems (zheng2025deepresearcher). In response, a novel class of Large Language Model (LLM)-based DR agents has emerged, systematically searching heterogeneous sources and synthesizing findings into information-dense reports (li2025webthinker).
Deep research agents, which autonomously conduct knowledge-intensive tasks by searching heterogeneous sources and synthesizing findings, are attracting significant attention in both academia (zhang2025web; xu2025comprehensive) and industry (openai-blog; gemini-dr; perplexity-dr; grok-dr). For example, OpenAI characterizes the functionality of such agents as the ability to produce comprehensive, human-quality reports by automating the often hours-long process of searching, analyzing, and synthesizing evidence (openai-blog). However, research in this area remains at an early stage. Specifically, comprehensive evaluation is crucial yet currently limited. Thus, in this study, we investigate DR and decompose it into three key aspects: Planning, Retrieval, and Reasoning. First, planning module decomposes a complex query into a well-structured sequence of sub-tasks. Second, retrieval module searches multiple sources, then iteratively selects useful evidence within a fixed budget. Third, reasoning module reasons over the defined sub-tasks, calls external search tools, and synthesizes them into a comprehensive report. To sum up, a comprehensive, modular-integrated evaluation paradigm can catalyze the development of stronger DR agents tailored to knowledge-intensive workloads. However, the corresponding benchmark remains nascent, with key bottlenecks outlined below.
Table 1: Comparison of Dr.Mi-Bench with existing DR benchmarks.
@@
Dr.Mi-Bench (Ours) Scientific
@@
Despite rapid progress, DR evaluation still exhibits two fundamental bottlenecks. First, prevailing benchmarks reduce the task to a retrieval slice, with little systematic evaluation of high-level planning and reasoning (wei2025browsecompsimplechallengingbenchmark; zhou2025browsecompzhbenchmarkingwebbrowsing; phan2025humanitysexam; mialon2023gaiabenchmarkgeneralai). Therefore, such evaluations do not faithfully reflect the capacity of an agent to solve complex problems. In particular, poor retrieval may be induced by an incorrect plan rather than shortcomings in the search process itself. Moreover, even when the correct evidence is retrieved, the model can still hallucinate, owing to limited semantic understanding and reasoning over long contexts, producing plausible yet incorrect conclusions. Second, current evaluations exhibit limited capability in scientific domains. existing DR benchmarks disproportionately emphasize general-purpose questions, leaving domain-specific and professional workloads, which is central to real DR use cases, underexplored (futuresearch2025deepresearchbenchevaluating; du2025deepresearchbenchcomprehensivebenchmark; java2025characterizingdeepresearchbenchmark).
@@
To address the aforementioned limitations, we formalize the workflow of DR agents and propose Dr.Mi-Bench, a modular-integrated deep research benchmark for scientific deep research agent. Notably, Dr.Mi-Eval evaluates the top-down capability of deep-research models to solve open-domain scientific research questions (Section 3). Our methodology is grounded in academic papers, which naturally provide structured guidelines for core competencies of an DR agent: planning, retrieval, and reasoning. For planning, we compare decomposed sub-tasks provided by agents against research plans annotated by experts based on the structural outline of papers, to quantify coverage of requisite sub-tasks, redundancy, and correctness. For retrieval, the benchmark assesses the relevance, coverage, and provenance of retrieved evidence, using the citations of papers as a verifiable ground truth. For reasoning, to evaluate reasoning on open-ended research questions, we construct a suite of diagnostics comprising boolean statements and answers based on the content of academic papers, which is used to evaluate the accuracy, consistency, depth, and breadth against the key findings. The three stages can be evaluated independently and can also be integrated into a closed loop to measure the end-to-end performance of a deep-research system. Furthermore, we evaluate the efficiency of DR agents by analyzing the trade-offs between their performance and two key costs: processing time (latency) and report length (token count). A detailed comparison between Dr.Mi-Bench and other existing DR benchmarks appears in Table 1. In support of this evaluation paradigm, we curate a human-annotated dataset spanning 10 scientific domains with 200 instances that include both research papers and review papers (Section 4). To demonstrate the utility of our benchmark, we conduct a comprehensive evaluation of both state-of-the-art end-to-end DR systems (e.g., OpenAI, Google) and a diverse set of foundational LLMs (e.g., GPT-o3, Llama) assessed as potential backbones. Our analysis reveals a fragmented performance landscape where agents exhibit specialized strengths: Grok in retrieval and Gemini in reasoning, yet all share common weaknesses, most notably in retrieving from complex review papers and maintaining consistent performance across diverse scientific domains. Dr.Mi-Bench reveals significant substantial capacity for reliable DR, enabling more trustworthy academic assistants.
@@
Figure 1: The Dr.Mi-Bench and Dr.Mi-Eval Framework. This figure illustrates how a DR agent workflow (a) is assessed by our modular-integrated evaluation paradigm (b) using our human-annotated benchmark (c). This approach enables a diagnostic analysis that pinpoints specific failures in the agent’s planning, retrieval, and reasoning capabilities.
@@ 2 Formalization of Deep Research Agent Workflow
Definition 1 (Deep Research Agent)
@@
$\pi:\ Q\!\to\!\mathcal{T},\quad\rho:\ (Q,\mathcal{T},\mathcal{E},\mathcal{H},\mathcal{B})\!\to\!\mathcal{E}^{+},\quad\sigma:\ (Q,\mathcal{T},\mathcal{E}^{+})\!\to\!\mathcal{R},$ (1)
where $\mathcal{T}=\langle T_{1},T_{2},\dots,T_{k}\rangle$ is a plan with ordered sub-tasks and $\mathcal{E}^{+}$ the curated set of predicted relevant evidence (citations). The detailed formalization and scoring for each component are provided in Appendix A.1 and Appendix A.15.
3 Dr.Mi-Bench Evaluation Paradigm
We introduce a novel modular-integrated evaluation paradigm for scientific DR agents, designed to assess their top-down research capabilities. Our evaluation is structured around three core modules: Planning, Retrieval, and Reasoning. It supports two complementary evaluation modes. In the End-to-end mode, modules are chained together for a holistic performance assessment, while in the isolated mode, each module is assessed independently with ground-truth inputs to measure its maximum capability. This dual-mode design combines the realism of a full-system evaluation with the precision of module-specific diagnostics, allowing us to attribute an agent’s performance to the intrinsic capabilities of its foundational LLM. This paradigm is implemented in our new benchmark, Dr.Mi-Bench, which is built upon a human-annotated dataset of 200 instances from 10 scientific domains. For comprehensive evaluation, each instance is annotated with a user query, a ground-truth plan, a set of gold evidence, and a suite of diagnostics.
Definition 2 (Modular-Integrated Evaluation Paradigm)
Our paradigm provides a dual-level assessment. We evaluate end-to-end agents using a End-to-end Mode for holistic performance, and foundational LLMs using an Isolated Mode for modular analysis of specific agentic capabilities. For each benchmark instance, we have a ground-truth tuple $(Q,\mathcal{T}^{\ast},\mathcal{E}^{+\ast},\mathcal{D})$ and corresponding metric suites ($\mathcal{M}_{\pi},\mathcal{M}_{\rho},\mathcal{M}_{\sigma}$).
$\textbf{End-to-end (Agents):}\quad\begin{rcases}\mathcal{T}\leftarrow\pi(Q)\\ \mathcal{E}^{+}\leftarrow\rho(Q,\mathcal{T},\mathcal{E},\mathcal{H},\mathcal{B})\\ R\leftarrow\sigma(Q,\mathcal{T},\mathcal{E}^{+})\end{rcases}$ $\implies\begin{dcases}\mathcal{M}_{\pi}(\mathcal{T},\mathcal{T}^{\ast})\\ \mathcal{M}_{\rho}(\mathcal{E}^{+},\mathcal{E}^{+\ast}).\\ \mathcal{M}_{\sigma}(R,\mathcal{D})\end{dcases}$ (2)
$\textbf{Isolated (LLMs):}\quad\begin{rcases}\mathcal{T}\leftarrow\pi(Q)\\ \mathcal{E}^{+}\leftarrow\rho(Q,\mathcal{T}^{\ast},\mathcal{E},\mathcal{H},\mathcal{B})\\ R\leftarrow\sigma(Q,\mathcal{T}^{\ast},\mathcal{E}^{+\ast})\end{rcases}$ $\implies\begin{dcases}\mathcal{M}_{\pi}(\mathcal{T},\mathcal{T}^{\ast})\\ \mathcal{M}_{\rho}(\mathcal{E}^{+},\mathcal{E}^{+\ast}).\\ \mathcal{M}_{\sigma}(R,\mathcal{D})\end{dcases}$ (3)
@@ 3.1 Planning Module Evaluation
The evaluation of the Planning module is designed to assess the ability of agents to decompose a query $Q$ into a structured plan. A plan predicted by agent, denoted as $\mathcal{T}=\langle T_{1},\dots,T_{k}\rangle$, is scored against the expert-annotated ground-truth plan, $\mathcal{T}^{\ast}=\langle T^{\ast}_{1},\dots,T^{\ast}_{m}\rangle$. Our protocol employs an exhaustive pairwise comparison between the predicted and ground-truth sub-tasks. Specifically, an LLM-based evaluator semantically compares each predicted sub-task $T_{i}\in\mathcal{T}$ against every ground-truth sub-task $T^{\ast}_{j}\in\mathcal{T}^{\ast}$ to determine relevance. This matrix of pairwise judgments is then used by our metric suite, $\mathcal{M}_{\pi}$, to quantify the plan’s overall coverage of essential sub-tasks, its structural correctness, and its redundancy. As planning is the initial stage of the pipeline, this evaluation process is identical for both the End-to-end and isolated modes.
@@ 3.2 Retrieval Module Evaluation
The evaluation of the retrieval module is designed to assess the consistency between the set of citations retrieved by the agent, $\mathcal{E}^{+}$, and the ground-truth citations annotated in our dataset, $\mathcal{E}^{+\ast}$. To quantify this alignment, our metric suite, $\mathcal{M}_{\rho}$, employs a rule-based exact match on canonical identifiers (DOIs) to calculate scores for relevance, coverage, and provenance (detailed in A.17).
We assess this capability in our two distinct modes: In the End-to-end Mode, the module receives the agent’s own predicted plan, $\mathcal{T}$, as input. This setup measures the retriever’s practical, end-to-end performance and reveals its susceptibility to any errors propagated from the upstream Planning module. In the Isolated Mode, the module is provided with the ground-truth plan, $\mathcal{T}^{\ast}$. This assesses the module’s maximum intrinsic capability, assuming a perfect plan was formulated. By comparing results from these two modes, we can effectively disentangle retrieval failures from planning failures.
@@ 3.3 Reasoning Module Evaluation
The reasoning module is evaluated on its core capability to synthesize a factually correct, coherent, and evidence-grounded scientific report $R$. Since report generation is an open-ended task that admits various valid solutions, direct comparison against a gold report can be unreliable. Therefore, we transform the subjective assessment of report quality into an objective and reproducible evaluation. Our diagnostic framework, $\mathcal{M}_{\sigma}$, scores the report $R$ against a set of predefined, ground-truth diagnostic pairs $\mathcal{D}$.
In the End-to-end Mode, the module receives the predicted plan $\mathcal{T}$ and predicted evidence set $\mathcal{E}^{+}$ from the preceding stages. This evaluation measures the final, overall quality of the agent’s generated report, capturing the cumulative impact of all potential upstream errors. In the Isolated Mode, the module is given the perfect, ground-truth plan $\mathcal{T}^{\ast}$ and evidence set $\mathcal{E}^{+\ast}$. This setup isolates the foundational LLM from relying on its inherent reasoning capability, revealing its performance when all prerequisite information is correct and complete. This dual-mode analysis allows us to distinguish true reasoning failures from performance degradation caused by propagated errors from the planning and retrieval modules.

### v2

# Beyond Retrieval: A Modular Benchmark for Academic Deep Research Agents

@@ Abstract
A surge in academic publications calls for automated deep research (DR) systems, but accurately evaluating them is still an open problem. First, existing benchmarks often focus narrowly on retrieval while neglecting high-level planning and reasoning. Second, existing benchmarks favor general domains over the academic domains that are the core application for DR agents. To address these gaps, we introduce ADRA-Bank, a modular benchmark for Academic DR Agents. Grounded in academic literature, our benchmark is a human-annotated dataset of 200 instances across 10 academic domains, including both research and review papers. Furthermore, we propose a modular Evaluation Paradigm for Academic DR Agents (ADRA-Eval), which leverages the rich structure of academic papers to assess the core capabilities of planning, retrieval, and reasoning. It employs two complementary modes: an end-to-end evaluation for DR agents and an isolated evaluation for foundational LLMs as potential backbones. Results reveal uneven capabilities: while agents show specialized strengths, they struggle with multi-source retrieval and cross-field consistency. Moreover, improving high-level planning capability is the crucial factor for unlocking the reasoning potential of foundational LLMs as backbones. By exposing these actionable failure modes, ADRA-Bank provides a diagnostic tool to guide the development of more reliable automatic academic research assistants.
Machine Learning, ICML
@@ 1 Introduction
Since human society has entered an era of unprecedented information explosion, conducting deep, accurate, and reliable research requires intensive knowledge work. Therefore, it is increasingly impractical for human experts to handle this task with conventional manual methods (Zhang et al., 2025a). The scale and speed of academic publication have reached unprecedented levels. For instance, more than 64 million research papers have become accessible online since 1996 (WordsRated, 2025), and the pace of new contributions continues to accelerate. The trend increases the difficulty of completing a comprehensive report with conventional manual methods, thereby underscoring the urgent need for automated systems (Zheng et al., 2025). In response, a novel class of Large Language Model (LLM)-based DR agents has emerged to conduct this complex workflow and produce information-dense reports automatically (Li et al., 2025b).
Deep research agents, which autonomously conduct knowledge-intensive tasks by searching heterogeneous sources and synthesizing findings, are attracting significant attention in both academia (Zhang et al., 2025b; Xu and Peng, 2025) and industry (OpenAI, 2025; Google, ; PerplexityAI, ; xAI, ). For example, OpenAI characterizes the functionality of such agents as the ability to produce comprehensive, human-quality reports by automating the often hours-long process of searching, analyzing, and synthesizing evidence (OpenAI, 2025). However, research in this area remains at an early stage. Specifically, comprehensive evaluation is crucial yet currently limited. Thus, in this study, we investigate DR and decompose it into three key aspects: Planning, Retrieval, and Reasoning. First, planning module decomposes a complex query into a well-structured sequence of sub-tasks. Second, retrieval module searches multiple sources, then iteratively selects useful evidence within a fixed budget. Third, reasoning module reasons over the defined sub-tasks, calls external search tools, and synthesizes them into a comprehensive report. In summary, a comprehensive, modular-integrated evaluation paradigm can accelerate the development of stronger DR agents designed for knowledge-intensive tasks. However, current benchmarks are still at an early stage.
Table 1: Comparison of ADRA-Bank with existing DR benchmarks.
@@
ADRA-Bank (Ours) Academic
@@
Despite rapid progress, DR evaluation still exhibits two fundamental bottlenecks. First, prevailing benchmarks reduce the task to a retrieval slice, with little systematic evaluation of high-level planning and reasoning (Wei et al., 2025; Zhou et al., 2025; Phan et al., 2025; Mialon et al., 2023). Therefore, such evaluations do not faithfully reflect the capacity of an agent to solve complex problems. In particular, poor retrieval may be induced by an incorrect plan rather than shortcomings in the search process itself. Moreover, even when the correct evidence is retrieved, the model can still hallucinate, owing to limited semantic understanding and reasoning over long contexts, producing plausible yet incorrect conclusions. Second, current evaluations demonstrate limited capability in academic domains. Existing DR benchmarks disproportionately emphasize general-purpose questions while neglecting domain-specific and professional workloads, which are central to real-world DR scenarios (FutureSearch et al., 2025; Du et al., 2025; Java et al., 2025).
@@
To address the aforementioned limitations, we formalize the workflow of DR agents and propose ADRA-Bank, a modular-integrated deep research benchmark for academic deep research agent. Notably, ADRA-Eval evaluates the top-down capability of deep-research models to solve open-domain academic research questions (Section 3). Our methodology is grounded in academic papers, which naturally provide structured guidelines for core competencies of an DR agent: planning, retrieval, and reasoning. For planning, we compare decomposed sub-tasks provided by agents against research plans annotated by experts based on the structural outline of papers, to quantify coverage of requisite sub-tasks, redundancy, and correctness. For retrieval, the benchmark assesses the relevance, coverage, and provenance of retrieved evidence, using the citations of papers as a verifiable ground truth. For reasoning, to evaluate reasoning on open-ended research questions, we construct a suite of diagnostics comprising boolean statements and answers based on the content of academic papers, which is used to evaluate the accuracy, consistency, depth, and breadth against the key findings. The three stages can be evaluated independently and can also be integrated into a closed loop to measure the end-to-end performance of a deep-research system. Furthermore, we evaluate the efficiency of DR agents by analyzing the trade-offs between their performance and three key costs: processing time (latency), report length (token count), and price (average and total). A detailed comparison between ADRA-Bank and other existing DR benchmarks appears in Table 1. In support of this evaluation paradigm, we curate a human-annotated dataset spanning 10 academic domains with 200 instances that include both research papers and review papers (Section 4). To demonstrate the utility of our benchmark, we conduct a comprehensive evaluation of both state-of-the-art end-to-end DR systems (e.g., OpenAI, Google) in Section 5 and a diverse set of foundational LLMs (e.g., GPT-o3, Llama) assessed as potential backbones in Section 6. Our analysis reveals a fragmented landscape: agents have distinct strengths (e.g., Grok in retrieval, Gemini in reasoning) but share critical struggles in retrieving from complex review papers and maintaining consistent performance across diverse academic domains. ADRA-Bank reveals significant substantial capacity for reliable DR agents, enabling development of trustworthy academic assistants.
@@
Figure 1: The ADRA-Bank and ADRA-Eval Framework. This figure illustrates how a DR agent workflow (a) is assessed by our modular-integrated evaluation paradigm (b) using our human-annotated benchmark (c). This approach enables a diagnostic analysis that pinpoints specific failures in the agent’s planning, retrieval, and reasoning capabilities.
@@ 2 Formalization of Deep Research Agent Workflow
Definition 2.1 (Deep Research Agent).
@@
$\pi$ $:Q\to\mathcal{T},$ (1)
$\rho$ $:(Q,\mathcal{T},\mathcal{E},\mathcal{H},\mathcal{B})\to\mathcal{E}^{+},$
$\sigma$ $:(Q,\mathcal{T},\mathcal{E}^{+})\to\mathcal{R},$
where $\mathcal{T}=\langle T_{1},T_{2},\dots,T_{k}\rangle$ is a plan with ordered sub-tasks and $\mathcal{E}^{+}$ the curated set of predicted relevant evidence (citations). The detailed formalization and scoring for each component are provided in Appendix A and Appendix F.
3 ADRA-Bank Evaluation Paradigm
We introduce a novel modular-integrated evaluation paradigm for academic DR agents, designed to assess their top-down research capabilities. Our evaluation is structured around three core modules: Planning, Retrieval, and Reasoning. It supports two complementary evaluation modes. In the end-to-end mode, modules are chained together for a holistic performance assessment, while in the isolated mode, each module is assessed independently with ground-truth inputs to measure its maximum capability. This dual-mode design combines the realism of a full-system evaluation with the precision of module-specific diagnostics, allowing us to attribute an agent’s performance to the intrinsic capabilities of its foundational LLM. This paradigm is implemented in our new benchmark, ADRA-Bank, which is built upon a human-annotated dataset of 200 instances from 10 academic domains. For comprehensive evaluation, each instance is annotated with a user query, a ground-truth plan, a set of gold evidence, and a suite of diagnostics.
Definition 3.1 (Modular-Integrated Evaluation Paradigm).
Our paradigm provides a dual-level assessment: end-to-end mode for holistic DR agent performance and isolated mode for modular LLM capabilities. For each benchmark instance, we have a ground-truth tuple $(Q,\mathcal{T}^{\ast},\mathcal{E}^{+\ast},\mathcal{D})$ and corresponding metric suites ($\mathcal{M}_{\pi},\mathcal{M}_{\rho},\mathcal{M}_{\sigma}$).
End-to-end (Agents):
$\begin{rcases}\mathcal{T}\leftarrow\pi(Q)\\ \mathcal{E}^{+}\leftarrow\rho(Q,\mathcal{T},\mathcal{E},\mathcal{H},\mathcal{B})\\ R\leftarrow\sigma(Q,\mathcal{T},\mathcal{E}^{+})\end{rcases}$ $\implies\begin{dcases}\mathcal{M}_{\pi}(\mathcal{T},\mathcal{T}^{\ast})\\ \mathcal{M}_{\rho}(\mathcal{E}^{+},\mathcal{E}^{+\ast})\\ \mathcal{M}_{\sigma}(R,\mathcal{D})\end{dcases}$ (3)
Isolated (LLMs):
$\begin{rcases}\mathcal{T}\leftarrow\pi(Q)\\ \mathcal{E}^{+}\leftarrow\rho(Q,\mathcal{T}^{\ast},\mathcal{E},\mathcal{H},\mathcal{B})\\ R\leftarrow\sigma(Q,\mathcal{T}^{\ast},\mathcal{E}^{+\ast})\end{rcases}$ $\implies\begin{dcases}\mathcal{M}_{\pi}(\mathcal{T},\mathcal{T}^{\ast})\\ \mathcal{M}_{\rho}(\mathcal{E}^{+},\mathcal{E}^{+\ast})\\ \mathcal{M}_{\sigma}(R,\mathcal{D})\end{dcases}$ (4)
@@ 3.1 Planning Module Evaluation
We evaluate the planning module by comparing the predicted plan $\mathcal{T}=\langle T_{1},\dots,T_{k}\rangle$ against the ground-truth $\mathcal{T}^{\ast}=\langle T^{\ast}_{1},\dots,T^{\ast}_{m}\rangle$. Our protocol employs an exhaustive pairwise comparison between the predicted and ground-truth sub-tasks. Specifically, an LLM-based evaluator semantically compares each predicted sub-task $T_{i}\in\mathcal{T}$ against every ground-truth sub-task $T^{\ast}_{j}\in\mathcal{T}^{\ast}$ to determine relevance. This matrix of pairwise judgments is then used by our metric suite, $\mathcal{M}_{\pi}$, to quantify the plan’s overall coverage of essential sub-tasks, its structural correctness, and its redundancy. As planning is the initial stage, this evaluation is identical for both the end-to-end and isolated modes.
@@ 3.2 Retrieval Module Evaluation
The evaluation of the retrieval module is designed to assess the consistency between the set of citations retrieved by the agent, $\mathcal{E}^{+}$, and the ground-truth citations annotated in our dataset, $\mathcal{E}^{+\ast}$. To quantify this alignment, our metric suite, $\mathcal{M}_{\rho}$, employs a rule-based exact match on canonical identifiers (DOIs) to calculate scores for relevance, coverage, and provenance (detailed in Appendix G).
We assess this capability in our two distinct modes: In the end-to-end mode, the module receives the agent’s own predicted plan, $\mathcal{T}$, as input. This setup measures the retriever’s practical, end-to-end performance and reveals its susceptibility to any errors propagated from the upstream Planning module. In the isolated mode, the module is provided with the ground-truth plan, $\mathcal{T}^{\ast}$. This assesses the module’s maximum intrinsic capability, assuming a perfect plan was formulated. Comparing results from two modes effectively disentangles retrieval failures from planning failures.
@@ 3.3 Reasoning Module Evaluation
The reasoning module is evaluated on its core capability to synthesize a factually correct, coherent, and evidence-grounded academic report $R$. Since report generation is an open-ended task that admits various valid solutions, direct comparison against a gold report can be unreliable. Therefore, we transform the subjective assessment of report quality into an objective and reproducible evaluation. Our diagnostic framework, $\mathcal{M}_{\sigma}$, scores the report $R$ against a set of predefined, ground-truth diagnostic pairs $\mathcal{D}$.
In the end-to-end mode, the module receives the predicted plan $\mathcal{T}$ and evidence set $\mathcal{E}^{+}$ from the preceding stages. This evaluation measures the final quality of the agent’s generated report, capturing the cumulative impact of all potential upstream errors. In the isolated mode, the module is given a ground-truth plan $\mathcal{T}^{\ast}$ and evidence set $\mathcal{E}^{+\ast}$. This setup isolates the foundational LLM from relying on its reasoning capability, revealing its performance when all prerequisite information is correct and complete. This dual-mode analysis distinguishes true reasoning failures from performance degradation caused by planning and retrieval modules.